# references.bib

### v1

@@ Abstract
In theatre, playwrights use the portrayal of characters to explore culturally based gender norms. In this paper, we develop quantitative methods to study gender depiction in the non-religious works (comedias) of Pedro Calderón de la Barca, a prolific Spanish 17th century author. We gather insights from a corpus of more than 100 plays by using a gender classifier and applying model explainability (attribution) methods to determine which text features are most influential in the model’s decision to classify speech as ‘male’ or ‘female’, indicating the most gendered elements of dialogue in Calderón’s comedias in a human accessible manner. We find that female and male characters are portrayed differently and can be identified by the gender prediction model at practically useful accuracies (up to f=0.83). Analysis reveals semantic aspects of gender portrayal, and demonstrates that the model is even useful in providing a relatively accurate scene-by-scene prediction of cross-dressing characters.
Keywords: automatic classification, Calderón de la Barca, Siglo de oro, machine learning
@@ 1 Introduction
In Spain, the Baroque period, was a period of immense artistic creativity, genereally known as the ”Golden Age” (siglo de oro). This is particularly true in literature, where the period saw exceptional writers such as Lope de Vega, Tirso de Molina or Pedro Calderón de la Barca. The latter, who lived from 1600 to 1681, is generally considered as as one of the most important playwrights of the age. He was immensely productive, writing a total of over 200 theatrical plays, both secular and religious, which had a lasting impact on Spanish theatre and beyond [fitzmaurice21]. He is particularly known for detailed and complex characterizations in his works [woods2013female].
@@
Not surprisingly, Calderón’s writings have been subject to intense analysis by literary scholars over a long period of time, and topics have moved in and out of fashion. For example, traditional foci of scholarship have been the role of honor and power in the works [greer2017play] or Calderón’s attention to dramatic structure [whitby1954structural].
A relatively new aspect among these is gender depiction, that is, the question of how Calderón conceptualized male and female roles in his plays differently, which has gained global attention in Hispanic Studies since the latter half of the 20th century ([arellano_ayuso_mujeres_2015, saez_veloz_2019, tietz_modo_2017]).
@@
Some studies indicate that Calderóns depicts a broadly period-typical understanding of gender identity and gender norms. For example, [ibarreche2020honor] study the representation of male characters, particularly in relation to honor in plays such as El médico de su honra, El Alcalde Zalamea, or El pintor de su deshonra. In these pieces, male characters, often husbands, fathers, or brothers, strive to uphold their social status, wielding power, owning properties, and navigating public spaces with traits such as bravery, patience, and caution [fernandez_guillermo_figura_2017]. Conversely, in plays like Casa con dos puertas mala es de guardar or La dama duende, women are frequently confined to domestic spaces, initiating disputes and deftly manipulating characters like puppets in games of seduction leading to a harmonious resolution [saez_veloz_2019].
@@
Meanwhile, other studies have found that Calderón’s work not only mirrors societal norms but also provides a dynamic platform for the exploration of gender identities [tietz_modo_2017]: ’Perverse’ women -women who use deception against men to achieve their goals - are predominantly present in tragic pieces, conspicuously absent in comedies [arellano_ayuso_mujeres_2015]. Furthermore, a significant number of plays depict female characters cross-dressing as men and performing traits typically associated with masculinity [bravo-villasante_mujer_1988, strosetzki_personajes_2017, araico2017lenguaje]. A commonly used trope of the time period, playwrights utilized cross-dressing as a means for female characters to transcend the domestic space and actively participate in both society and the action of the play. The aforementioned scholarship draws attention conflicting tendencies of gender depiction in Calderón’s work: it is both deeply shaped by plot and genre conventions, but is used as a grounds to expand social ideas about gender roles by representing women with agency who drive plot progression and utilize covert ways to exercise their will [woods2013female]. A possible explanation for the picture that emerges is that, throughout the massive catalogue of Calderón’s works, a select few have been given close attention by scholars who adopt a ’close reading’ method focusing on individual dramas sampled of the more than 100 secular dramas written by Calderón. Consequently, studying the all dramas can show if this picture holds true for the oeuvre as a whole.
@@
This prompts the classical motivation for scalable reading methods [weitin17] from digital humanities, combining computational models with manual interpretation: Can we make out global patterns in gender depiction that hold across Calderón’s whose body of secular dramas?
@@
More specifically, we propose to build on recent developments regarding the explainability of AI models [wiegreffe-pinter-2019-attention, bibal-etal-2022-attention] which aim at providing a human-understandable justification of model decisions. This makes it possible to consider a computational model of a literary phenomenon primarily as an aggregation mechanism that summarizes large amounts of text into a much smaller set of predictions, and then to analyze these predictions back to the linguistic properties that gave rise to them.
@@
In this study, we show the feasibility of this research program by training a gender classifier based on the characters’ speech. The motivation here is not to carry out the task of gender classification, but to ask questions about the classifier’s predictions: How difficult is the task — i.e., how consistent is Calderón’s gender depiction? What features does the model rely on for its predictions — i.e., what linguistic patterns does Calderón use for gender depiction? What characters or character groups are easy or difficult to classify — i.e., what characters are depicted in unconventional or unexpected fashions?
@@
Our work is not the first to classify characters and their properties based on their speech. For instance, [vishnubhotla-etal-2019-fictional] analyze the ability of classifiers to distinguish among characters within a play to determine if authors were able to write characters with distinct voices. The closest precursor of our study is [bullard-ovesdotter-alm-2014-computational] who predict various demographic properties of characters, including gender. However, their study was still based on hand-picked features fed into a regression model. [savoy2023stylometric] also conducted a classification experiment on the gender of Shakespeare characters based on lines, using random forest and logistic regression.
@@
Instead, we work with the more recent language models (LMs) which give rise to two major new design questions. First, LMs promise to do away with feature engineering, but for the purpose we have in mind (explaining the model predictions), not all parts of the input are equally informative. Second, LMs are still known to struggle with identifying relevant information in long texts. This raises the question of how much information to present to the classifier at each time: Too little (e.g., just one line), and the input may contain insufficient hints as to the speaker’s gender; too much (e.g., a character’s complete text from a drama), and the model may struggle. Also, the predictions of the model for each individual part of the input have to be aggregated into a prediction at a more global level, for which we also consider multiple options. We discuss both of these issues in Section 2 below.
@@
Our models are able to correctly classify the gender of characters in Calderón’s comedias. Our experiments yield the best results given the most amount of input text, but the models also perform well using an aggregation method that takes into account the confidence of each prediction. We take this as a positive result for computational analysis of gender portrayal in the works of Calderón: it shows that there are recognizable gender cues at the scene level and, at the same time that, characters’ gender roles can vary over the course of a play. Finally, an examination of the most explanatory features in the models predictions demonstrate that the model picks up on both grammatical and semantic information that contribute to gender prediction. The examination of cross-dressing female character demonstrates that these character’s lines are more similar to the speech of male characters compared to more traditional female characters, particularly during cross-dressing scenes.
@@ 2 Methods
Language models and word embedding-based classification.
@@
In the first phase, distributional semantics, built on ideas from structuralist semantics [Firth1957], was developed into an approach that aimed at capturing the meaning of target words by counting co-occurrences of these target words with words in their context. This naturally corresponds to a representations of words as context vectors (or embeddings), and of semantic similarity as a measure of contextual similarity between words [miller91:contextual].
@@
In the second phase, the construction of such embeddings moved from explicit context counting to the use of neural language models. Such language models are presented with a textual prefix (“holiday on Sri …”) and have the task of predicting the subsequent token (“Lanka”). When trained to do so on large text corpora, the learned parameters of such language models are essentially word embeddings [Mikolov:2013:DRW:2999792.2999959].
@@
The next phase involved the introduction of the so-called transformer architecture, which introduces to ability to make word embeddings sensitive on their context: e.g., “root” receives a different embedding in the phrase “root canal” than in the phrase “square root”. The resulting contextualized language models can account for lexical ambiguity in a principled fashion [devlin19:_bert]. Model families of this type, such as BERT, have shown to be applicable to almost all standard tasks in NLP by obtaining word embeddings for an input text, aggregating these embeddings to obtain a text embedding (e.g. simply by averaging), and passing the aggregate through a very simple decision mechanism – e.g., a single-layer neural network with two output units which directly corresponds to a logistic regression model over the text embedding. Such models are typically trained in two phases: first the language model is ’pre-trained’ on a large amount of unlabeled text, then the decision layer is ’fine-tuned’ on a comparatively small amount of task-specific labeled text.
@@
The final phase in the development of language models consists in the emergence of large language models (LLMs) which result from massive scaling up the phase-3 models. As a result, LLMs are to a substantial degree task-agnostic (i.e., do not require fine-tuning) and can be instructed (“prompted”) directly to carry out some task [10.1145/3560815]. To solve a classification task, for example, models can be asked to answer ’yes’ or ’no’ similar to the way a human annotator would be instructed.
@@
For the task we consider in this study – predicting the gender of a character in a drama based on (some part of) their utterances – both contextualized embedding models and LLMs are plausible candidates to develop classifiers. We decided to build our classifier on top of a contextualized embedding model. Such models are typically much smaller, (and thus computation efficient), they are more transparent (the training process of most LLMs is confidential [10.1145/3571884.3604316]), and prompt-based approaches face problems regarding robustness and bias [ceronetal24].
Design decisions for gender prediction: Input granularity and aggregation.
@@
As outlined in the Introduction, a central question that must be decided when presenting text to a LM for classification is the length of the input which we will call input granularity. Short inputs may consist of utterances which are not informative regarding the speaker’s gender. As for long inputs, older LM-based models have generally a limited input size (often 512 tokens [^1] See Section 3 for details about tokenization.), and while newer LLM-based models can deal with much longer inputs, they struggle with identifying the relevant information, in particular if it in the middle of the text [liu23-lostinthemiddle].
@@
Figure 1: Gender classification at different levels of granularity (blue = masculine, red = feminine) This image represents experimental setup method employed in this paper.
@@
Figure 1 shows a simplified view of the hierarchical structure of a character’s utterances in a drama: At the bottom we have individual utterances, which vary in length from single sentences to long monologues. Utterances can be grouped together at the scene level, and then again at the global / character level. Since we do not know a priori which of these levels works best, our experimental setup will evaluate models that have been trained at each of these three levels. In Figure 1, male predictions are shown in blue, and female predictions in red.
@@
If we choose to classify at a more fine-grained level (such as utterances), there is no guarantee that all utterances of one character are classified in the same way – in fact, given our assumption that some at least some utterances are uninformative regarding gender, complete consistency would be surprising (cf. Utterances 2 vs. 3 in Figure 1). Therefore, we also need to define an aggregation method to return a character-level response based on the concrete utterance- or scene-level predictions. [^2] Utterance-level predictions could also be aggregated at the scene level. We do not considerthis option to simplify the experimental setup. We consider two aggregation strategies. The simple one is majority vote: we count the number of utterances (or scenes) predicted as male or female, respectively, and assign the more frequent gender. In Figure 1, 3 out of 4 utterances are classified as male, which leads to an overall classification of the character as male. The second strategy makes use of the fact that LM-based models do not just output labels, but also probabilities. Assuming that more informative inputs are classified more confidently, taking these probabilities into account should give more weight to informative utterances. Formally, we use the geometric mean $p_{GM}(g|c)$ of a gender $g$ for a character $c$ given the individual inputs $i_{c}$ for the character:
$p_{GM}(g|c)=\left(\prod_{i_{c}=1}^{x}p(g|i_{c})\right)^{\frac{1}{x}}$ (1)
@@
In Figure 1, the saturation of the colors indicates confidence: In scene 1, where the character speaks three times, utterance 1 is classified as male with a low probability, utterance 2 is classified as male with a high probability, and utterance 3 is classified as female with a low probability, leading to a character-level aggregated classification of male.
Analyzing neural classifiers with attribution methods.
The topic of making machine learning models explainable has seen considerable attention in recent years, from a variety of motivations [https://doi.org/10.1002/hbe2.117]. Two major families [Murdoch] are model-based approaches – which build on the inherent interpretability of small models – and post-hoc methods which analyze previously trained models to explain their predictions. In turn, post-hoc methods can be grouped into global and local methods. Global methods analyze the entirety of a model’s predictions over some dataset, while local models explain individual predictions [atanasova-etal-2020-diagnostic]. In our study, we focus on local methods because we are interested in looking at individual predictions. A prominent local approach for the attention-based models ubiquitous in NLP is the analysis of attention weights; however, it has been shown that these models are not guaranteed to faithfully attribute the model’s prediction to its input features [jain-wallace-2019-attention]. Therefore, we adopt the so-called integrated gradient approach by [sundararajan2017axiomaticattributiondeepnetworks]. Applied to a binary classification task with classes A and B, integrated gradients explains every prediction as a sum of weights (positive, i.e, supporting class A, or negative, i.e., supportcing class B) over the input features. In our case, every word in the input text is assigned a weight that indicates how strongly is supports (or contradicts) the assignment of a particular gender to the text. Figure 2 shows some example sentences from a Calderón drama, where orange color indicates female speech and blue color male speech, and where saturation corresponds to the strength of the cue.
@@
Lexical overfitting.
A final complication in our use of explanation methods to mine a classifier for information about an author’s linguistic pattern is that the features presented to the classifier need to be informative for the research question. For example, person names are arguably uninformative for gender portrayal: the fact that men and women interact more within their genders than across is not an interesting insight. More generally, the problem is the Zipf distribution of natural language, as a consequence of which there is a large number of infrequent words that occur either only in male or in female speech and could therefore be identified by the classifier as highly informative for the respective class. In Calderón’s tragedy Un castigo en tres venganzas, for example, only the main character Federico ever addresses the character Becoquín – meaning when the token Becoquín appears, it is with certainty spoken by the character Federico and thus a perfect cue for male gender.
This can be seen as an instance of overfitting [wolfe-caliskan-2021-low]. The problem is also well known in stylometry [10.1093/llc/fqx023]. It is usually addressed by restricting attention to the top-$n$ words or some mid-frequent range of words [caliskan2022gender]. Another strategy is to mask particularly predictive tokens [tang2022alleviating], thus making them unavailable to the classifier to forcing it to focus on more general cues.

### v2

# Towards Computational Analysis of Gender Depiction in the Comedias of Calderón de la Barca

@@ Abstract
In theater, playwrights can use the portrayal of characters to explore culturally based gender norms. In this paper, we develop quantitative methods to study gender depiction in the non-religious works (comedias) of Pedro Calderón de la Barca, a prolific Spanish 17th century author. We gather insights from a corpus of more than 100 plays by using a gender classifier and applying model explainability (attribution) methods to determine which text features are most influential in the model’s decision to classify speech as ‘male’ or ‘female’, indicating the most gendered elements of dialogue in Calderón’s comedias in a manner accessible to humans. We find that the classifier picks up the male/female distinction mostly successfully (up to f=0.83), yields semantic aspects of gender portrayal, and demonstrates that the model is useful in providing a relatively accurate scene-by-scene prediction of cross-dressing characters.
@@ 1 Introduction
In Spain, the Baroque period (approx. 1600–1700) was a period of immense artistic creativity, generally known as the ‘Golden Age’ (Siglo de Oro). This is particularly true in literature, where the period saw exceptional writers such as Lope de Vega, Tirso de Molina and Pedro Calderón de la Barca. The latter, who lived from 1600 to 1681, is generally considered as one of the most important playwrights of the age. He was immensely productive, writing a total of over 200 theatrical plays, both secular and religious, which had a lasting impact on Spanish theatre and beyond ([fitzmaurice_1921]). He is particularly known for detailed and nuanced portrayal of the characters in his works ([woods_2013]).
@@
Not surprisingly, Calderón’s writings have been subject to intense analysis by literary scholars over a long period of time, and topics have moved in and out of fashion. For example, traditional foci of scholarship have been the role of honor and power ([greer_2017]) as well as his attention to dramatic structure ([whitby_1954]).
@@
A relatively new aspect among these, which has gained global attention in Hispanic Studies since the latter half of the 20th century, is gender depiction, that is, the question of how Calderón conceptualized male and female roles in his plays differently ([arellano_2015, saez_2019, tietz_2017]). Interestingly, scholars have come to very different conclusions regarding Calderón’s gender portrayal. Some studies find that Calderón depicts a broadly period-typical understanding of gender identity and gender norms. For example, [ibarreche_2020] study the representation of male characters, particularly in relation to honor, in plays such as El médico de su honra, El Alcalde Zalamea, or El pintor de su deshonra. In these pieces, male characters, often husbands, fathers, or brothers, strive to uphold their social status, wielding power, owning properties, and navigating public spaces with traits such as bravery, patience, and caution ([fernandez_2017]). Conversely, in plays like Casa con dos puertas mala es de guardar or La dama duende, women are frequently confined to domestic spaces, initiating disputes and deftly manipulating characters like puppets in games of seduction leading to a harmonious resolution ([saez_2019]).
In contrast, other studies have found that Calderón’s work transcends societal norms by providing a dynamic platform for the exploration of gender identities ([tietz_2017]) such as “perverse” women – women who use deception against men to achieve their goals. While absent from comedies, these form a recurring character type in tragic pieces ([arellano_2015]). Similarly, several well-known works depict female characters cross-dressing as men and performing traits typically associated with masculinity ([bravo_1976, strosetzki_2017, araico_2017]). A commonly used trope of the time period, playwrights utilized cross-dressing as a means for female characters to transcend the domestic space and actively participate in both society and the action of the play.
@@
Thus, there appear to be conflicting tendencies of gender depiction in Calderón’s work: It is both deeply shaped by plot and genre conventions, but is used as a grounds to expand social ideas about gender roles by representing women with agency who drive plot progression and utilize covert ways to exercise their will ([woods_2013]). A possible explanation for the picture that emerges is that, throughout the massive catalogue of Calderón’s works, a select few have been given close attention by scholars who adopt a ‘close reading’ method focusing on individual dramas sampled of the more than 100 secular dramas written by Calderón. Consequently, studying all the dramas can show if this picture holds true for the oeuvre as a whole.
@@
This prompts the classical motivation for scalable reading methods ([weitin_2017]) from digital humanities, combining computational models with manual interpretation: Can we make out global patterns in gender depiction that hold across Calderón’s whole body of secular dramas?
@@
In this study, we propose to build on recent developments regarding the explainability of AI models ([wiegreffe_2019, bibal_2022]) which aim at providing a human-understandable justification of model decisions. This makes it possible to consider a computational model of a literary phenomenon primarily as an aggregation mechanism that summarizes large amounts of text into a much smaller set of predictions, and then to trace these predictions back to the linguistic properties that gave rise to them.
@@
We show the feasibility of this research program by training a gender classifier based on the characters’ speech. The motivation here is not to carry out the task of gender classification, but to ask questions about the classifier’s predictions: How difficult is the task – i.e., how consistent is Calderón’s gender depiction? What features does the model rely on for its predictions – i.e., what linguistic patterns does Calderón use for gender depiction? What characters or character groups are easy or difficult to classify – i.e., what characters are depicted in unconventional or unexpected fashions?
@@
Our work is not the first to classify characters and their properties based on their speech. Five are particularly close. First, [vishnubhotla_2019] analyze the ability of classifiers to distinguish among characters within a play to determine if authors were able to write characters with distinct voices; they however ask the question at a general level, without reference to gender or Calderón. Second, [bullard_2014] predict various demographic properties of characters, including gender. However, their study was still based on hand-picked features fed into a regression model. Next, [savoy_2023] conducted a classification experiment on the gender of Shakespeare characters based on lines, using random forest and logistic regression. Two more studies adopt tools from stylometry, extending the approach from authorship attribution to the more fine-grained character portrayal. [vsecla_2024] carry out $n$-gram analysis to examine characteristics of gender depiction in French, Russian, German, and English dramas, specifically looking at the distinctiveness of characters and revealing distinct authorial style. Finally, [lorenzo_2024] compares the depiction of male and female characters in the works of Calderón and another playwright, Sor Juana de la Cruz, on a comparatively small number of works.
@@
Our study takes advantage of recent developments in Natural Language Processing by representing character speech using language models (LMs). This gives rise to two major new design questions. First, LMs decrease reliance on feature engineering, but for the purpose we have in mind (explaining the model predictions), not all parts of the input are equally informative. Second, LMs are still known to struggle with identifying relevant information in long texts. This raises the question of how much information to present to the classifier at each time: Too little (e.g., just one line), and the input may contain insufficient hints as to the speaker’s gender; too much (e.g., a character’s complete text from a drama), and the model may struggle. Also, the predictions of the model for each individual part of the input have to be aggregated into a prediction at a more global level, for which we also consider multiple options. We discuss both of these issues in section 2 below.
@@
We find that our models are able to correctly classify the gender of characters in Calderón’s comedias. Our experiments yield the best results when given the most amount of input text, but the models also perform well when using an aggregation method that takes into account the confidence of each prediction. We take this as a positive result for computational analysis of gender portrayal in the works of Calderón: It shows that there are recognizable gender cues at the scene level and, at the same time, that characters’ gender roles can vary over the course of a play. Finally, an examination of the most explanatory features in the models predictions demonstrate that the model picks up on both grammatical and semantic information that contribute to gender prediction. The examination of cross-dressing female characters demonstrates that these characters’ lines are more similar to the speech of male characters compared to more traditional female characters, particularly during cross-dressing scenes.
@@ 2 Methods
2.1 Language Models and Word Embedding-Based Classification
@@
In the first phase, distributional semantics, built on ideas from structuralist semantics ([firth_1957]), was developed into an approach that aimed at capturing the meaning of target words by counting co-occurrences of these target words with words in their context. This naturally corresponds to a representation of words as context vectors (or embeddings), and of semantic similarity as a measure of contextual similarity between words ([miller_1991]).
@@
In the second phase, the construction of such embeddings moved from explicit context counting to the use of neural language models. Such language models are presented with a textual prefix (‘holiday on Sri …’) and have the task of predicting the subsequent token (‘Lanka’). When trained to do so on large text corpora, the learned parameters of such language models are essentially word embeddings ([mikolov_2013]).
@@
The next phase involved the introduction of the so-called transformer architecture, which introduces the ability to make word embeddings sensitive to their context: e.g., ‘root’ receives a different embedding in the phrase ‘root canal’ than in the phrase ‘square root’. The resulting contextualized language models can account for lexical ambiguity in a principled fashion ([devlin_2019]). Model families of this type, such as BERT, have shown to be applicable to almost all standard tasks in NLP by obtaining word embeddings for an input text, aggregating these embeddings to obtain a text embedding (e.g. simply by averaging), and passing the aggregate through a very simple decision mechanism – e.g., a single-layer neural network with two output units which directly corresponds to a logistic regression model over the text embedding. Such models are typically trained in two phases: First the language model is ‘pre-trained’ on a large amount of unlabeled text, then the decision layer is ‘fine-tuned’ on a comparatively small amount of task-specific labeled text.
@@
The final phase in the development of language models consists in the emergence of large language models (LLMs) which result from massive scaling up of the model training. The resulting LLMs are to a substantial degree task-agnostic (i.e., do not require fine-tuning) and can be instructed (‘prompted’) directly to carry out some task ([liu_2023a]). To solve a classification task, for example, models can be asked to answer ‘yes’ or ‘no’ similar to the way a human annotator would be instructed.
@@
For the task we consider in this study – predicting the gender of a character in a drama based on (some part of) their utterances – both contextualized embedding models and LLMs are plausible candidates to develop classifiers. We decided to build our classifier on top of a contextualized embedding model. Such models are typically much smaller (and thus computation efficient), they are more transparent (the details of the training processes of almost all LLMs are confidential ([liesenfeld_2023])), and prompt-based approaches face problems regarding robustness and bias ([webson_2022]).
2.2 Design Decisions: Input Granularity and Aggregation
@@
As outlined above, a central question that must be decided when presenting text to a LM for classification is the length of the input for which the model is supposed to make a single prediction, that is, the input granularity. Short inputs may consist of utterances which are not informative regarding the speaker’s gender. As for long inputs, older LM-based models have generally a limited input size (often 512 tokens) [^1] See section 3 for details about tokenization. and while newer LLM-based models can deal with much longer inputs, they struggle with identifying the relevant information, in particular if it is in the middle of the text ([liu_2023b]).
@@
Figure 1: Gender classification for a character at different levels of granularity (blue = masculine, red = feminine). Saturation indicates model confidence.
@@
Figure 1 shows a simplified view of the hierarchical structure of a character’s utterances in a drama: At the bottom we have individual utterances, which vary in length from single sentences to long monologues. Utterances can be grouped together at the scene level, and then again at the global character level. Models can be trained to make gender predictions at each of these levels (shown in blue and red, respectively), and predictions can vary in confidence (shown by saturation). Since we do not know a priori which of these levels works best, our experiment will evaluate models that have been trained at each of these three levels.
@@
If we choose to classify at a more fine-grained level (such as utterances), there is no guarantee that all utterances of one character are classified in the same way – in fact, given our assumption that at least some utterances are uninformative regarding gender, complete consistency would be surprising (utterances 2 vs. 3 in Figure 1). Therefore, we also need to define an aggregation method to return a character-level response based on the concrete utterance- or scene-level predictions. [^2] Utterance-level predictions could also be aggregated at the scene level. We do not consider this option to simplify the experimental setup. We consider two aggregation strategies. The simple one is majority vote: We count the number of utterances (or scenes) predicted as male or female, respectively, and assign the more frequent gender. In Figure 1, 3 out of 4 utterances are classified as male, which leads to an overall classification of the character as male. The second strategy makes use of the fact that LM-based models do not just output labels, but also probabilities. Assuming that more informative inputs are classified more confidently, taking these probabilities into account should give more weight to informative utterances. Formally, we use the geometric mean $p_{GM}(g|c)$ of a gender $g$ for a character $c$ given the individual utterances $u(c)$ for the character:
$p_{GM}(g|c)=\left(\prod_{u(c)\in U(c)}^{x}p(g|i_{c})\right)^{\frac{1}{x}}$ (1)
@@
In Figure 1, utterance 1 is low probability male, utterances 2 and 4 are high probability male, and utterance 3 is low probability female, leading to a character-level aggregated classification of male.
@@
2.3 Analyzing Neural Classifiers with Attribution Methods
The topic of making machine learning models explainable has seen considerable attention in recent years, from a variety of motivations ([riedl_2019]). Two major families ([murdoch_2019]) are model-based approaches – which build on the inherent interpretability of small models – and post-hoc methods – which analyze previously trained models to explain their predictions. In turn, post-hoc methods can be grouped into global and local methods. Global methods analyze the entirety of a model’s predictions over some dataset, while local models explain individual predictions ([atanasova_2020]). In our study, we focus on local methods because we are interested in looking at individual predictions. A prominent local approach for the attention-based models ubiquitous in NLP is the analysis of attention weights; however, it has been shown that these models are not guaranteed to faithfully attribute the model’s prediction to its input features ([jain_2019]). Therefore, we adopt the so-called integrated gradient approach by [sundararajan_2017]. Applied to a binary classification task with classes A and B, this approach explains every prediction as a sum of weights (positive, i.e, supporting class A, or negative, i.e., supporting class B) over the input features. In our case, every word in the input text is assigned a weight that indicates how strongly it supports (or contradicts) the assignment of a particular gender to the text.
Figure 2 shows some example sentences from a Calderón drama, where orange color indicates words that are typical for female speech and blue color for male speech. Again, saturation corresponds to the strength of the cue. Note that these attributions can combine morphological and semantic considerations: The most strongly female token, “desesperada” (desperate), is more likely to be uttered by a woman both due to its female suffix -a, and by topical association.
2.4 Lexical Overfitting
A final complication in our use of explanation methods to mine a classifier for information about an author’s linguistic patterns is that the features presented to the classifier need to be informative for the research question. This is not guaranteed, since a classifier can base its decision on features that are uninteresting from a theoretical perspective. For example, person names are informative for gender classification since men and women interact more within their genders than across. At the same time, person names are arguably uninteresting with respect to gender portrayal. In Calderón’s tragedy Un castigo en tres venganzas, for example, only the main character Federico ever addresses the character Becoquín – meaning when the token Becoquín appears, it is with certainty spoken by the character Federico and thus a perfect cue for male gender. More generally, the problem is the Zipf distribution of natural language, as a consequence of which there is a large number of infrequent words that occur either only in male or in female speech and could therefore be identified by the classifier as highly informative for the respective class.
This can be seen as an instance of overfitting ([wolfe_2021]). The problem is also well known in stylometry ([evert_2017, vsecla_2024]). It is usually addressed by requiring words to meet certain frequency conditions, such as top-$n$ words, mid-frequent words ([caliskan_2022]) or words that occur in a minimum number of documents. Another strategy is to mask particularly predictive tokens ([tang_2022]), thus making them unavailable to the classifier and forcing it to focus on more general cues.